A and B are independent events. Find the missing probability.
P(B) = 1/5, P(A and B) = 2/13, Find P(A)

Answer by stanbon(75887)

A and B are independent events. Find the missing probability.
P(B) = 1/5, P(A and B) = 2/13, Find P(A)
------------
P(A and B)
= P(A)*P(B) = 2/13
= P(A)*(1/5) = 2/13
P(A) = (2/13)/(1/5) = (2/13)*(5/1) = 10/13
=============================================
Cheers,
Stan H.